# Near-Optimal Decision Trees in a SPLIT Second

### v2

@@ A.9 Greedy Algorithm
0:  $D,d,\lambda$Data subset, depth constraint, leaf regularization}
0:  $t_{\textrm{greedy}},lb$tree grown with a greedy, CART-style method; and the objective of that tree}
1:  $t_{\textrm{greedy}}\leftarrow$ (Leaf predicting the majority label in $D$)
@@ A.10 RESPLIT Algorithm
Algorithm 5 $\ell$, D, $\lambda$, $d_{l}$, $d$)
0:  $\ell$, $D$, $\lambda$, $d_{l}$, $d$loss function, samples, regularizer, lookahead depth, depth budget}
1:  ModifiedTreeFARMS = TreeFARMS reconfigured to use get__\__bounds (Algorithm 1) whenever it encounters a new subproblem
1:  $tf=$ ModifiedTreeFARMS$(\ell,D,\lambda,d_{l})$Call ModifiedTreeFARMS with depth budget $d_{l}$}
@@
7:        $T_{g},L_{g}=$ Greedy($D(u)$,$d-d_{u}$, $\lambda_{u}$) {Objective of greedy tree trained on subproblem}
8:        $t_{u}=$ TreeFARMS$(D(u),d-d_{u},\lambda_{u},L_{g})$Find all subtrees with loss less than $L_{g}$}
@@
13:     $t_{lookahead}=$ Enumerate__\__TreeFARMS__\__subtrees {For each node in this prefix tree, store the number of subtrees we can generate rooted at that node. This speeds up indexing}
@@
15:  return  $tf$Return in-place edited ModifiedTreeFARMS object}
@@ A.11 Indexing Trees in RESPLIT
  • Once the cumulative count is known, we start looping over the entire Rashomon set. For the $i^{th}$ index, we first obtain the corresponding prefix tree and then find the relative index of the $i^{th}$ tree within this prefix tree structure. For example, if we query the $500^{th}$ tree and our prefix contains trees indexed $400-600$ in the Rashomon set, the relative index of the query tree within this prefix is $100$.
  • Using Algorithm 8, we proceed to recursively locate the relevant subtrees beyond the prefix. In particular, at a given node in the prefix, we have access to the number of sub-trees that can be formed with its left and right children. We use this information to create two separate indexes for the left and right child (seen in lines $9-10$)
@@
Algorithm 6 Enumerate__\__TreeFARMS__\__subtrees
0:  $t_{lookahead}$Lookahed prefix with TreeFARMS objects attached to leaves}
@@
4:     Return $len$($t_{lookahead}$), $t_{lookahead}$
@@
6:  left__\__expansions, left__\__subtree $=$ enumerate__\__treefarms_subtrees($t_{lookahead}$.left__\__child)
7:  $t_{lookahead}$.left__\__child.node $=$ left__\__subtree
8:  $t_{lookahead}$.left__\__child.subtree__\__count $=$ left_expansions
9:  right__\__expansions, right__\__subtree $=$ enumerate__\__treefarms__\__subtrees($t_{lookahead}$.right__\__child)
10:  $t_{lookahead}$.right__\__child.node $=$ right__\__subtree
11:  $t_{lookahead}$.right__\__child.subtree__\__count $=$ right_expansions
12:  Return left__\__expansions $\times$ right__\__expansions, $t_{lookahead}$Total number of subtrees = cross product of left and right subtree count}
Algorithm 7 RESPLIT__\__Rset__\__Count(RESPLIT__\__obj)
0:  RESPLIT__\__obj {The RESPLIT object output by Algorithm 5}
1:  $t_{count}=0$Total ##\## trees}
2:  $p_{counts}=[]$Cumulative count of ##\## trees beginning with a given prefix}
3:  for $t_{lookahead}\in$ RESPLIT__\__obj do
@@
7:        $s_{count}=len(tf_{u})$Number of subtrees found for subproblem $D(u)$}
@@
Algorithm 8 get_leaf_subtree_at_idx($t_{lookahead}$, tree__\__idx)
0:  $t_{lookahead}$, tree__\__idx {A lookahead prefix tree with TreeFARMS objects attached to leaves, index to search within this tree}
@@
6:  $tree\leftarrow$ Node($t_{lookahead}$.feature)) {Initialize an empty node}
7:  left__\__count $=t_{lookahead}$.left__\__child.subtree__\__count {The number of subtrees that can be found rooted at this node}
8:  right__\__count $=t_{lookahead}$.right_child.subtree__\__count
9:  right__\__idx $=$ tree__\__idx % right__\__count
10:  left__\__idx $=$ tree__\__idx // right__\__count
11:  $tree$.left__\__child $=$ get_leaf_subtree_at_idx($t_{lookahead}$.left_child.node, left__\__idx)
12:  $tree$.right__\__child $=$ get_leaf_subtree_at_idx($t_{lookahead}$.right_child.node, right__\__idx)
13:  return  $tree$
Algorithm 9 RESPLIT__\__indexing
@@
2:  $t_{count},p_{counts}$ = RESPLIT_Rset_Count(RESPLIT_obj) {Total number of trees and prefix-wise cumulative counts}
@@
6:        start = p_counts[$i-1$]+1 {Start index for prefix $i$}
@@
8:     end = p_counts[$i$] {End index for prefix $i$}
@@ A.12 Modifications to Existing GOSDT / TreeFARMS Code
In this section, we detail the main modifications we made to the existing GOSDT and TreeFARMS codebase in order to set up SPLIT, LicketySPLIT, and RESPLIT. The algorithm components in red are the modifications - note that GOSDT and TreeFARMS both call these functions. TreeFARMS does some additional post-processing of the search trie to find the set of near optimal trees - the details can be seen in (Xin et al., 2022).
Algorithm 10 $\ell$, $D$, $\lambda$, $d_{l}$, $d$)
0:  $\ell$, $D$, $\lambda$, $d_{l}$, $d$loss function, dataset, regularizer, lookahead depth, global depth budget}
1:  $Q\leftarrow\emptyset$ priority queue}
2:  $G\leftarrow\emptyset$ dependency graph}
3:  $s_{0}\leftarrow\{1,\ldots,1\}$ bit-vector of 1’s of length $n$}
4:  $p_{0}\leftarrow\text{FIND\_OR\_CREATE\_NODE}(G,s_{0},{\color[rgb]{1,0,0}d_{l}},{\color[rgb]{1,0,0}d},{\color[rgb]{1,0,0}0})$ root (with depth 0)}
5:  $Q.\text{push}((s_{0},{\color[rgb]{1,0,0}0}))$ add to priority queue}
6:  while $p_{0}.\text{lb}\neq p_{0}.\text{ub}$ do
7:     $s,{\color[rgb]{1,0,0}d}^{\prime}\leftarrow Q.\text{pop}()$index of problem to work on}
8:     $p\leftarrow G.\text{find}(s)$ find problem to work on}
9:     if $p.\text{lb}=p.\text{ub}$ then
10:        continue { problem already solved}
11:     end if
12:     $(lb^{\prime},ub^{\prime})\leftarrow(\infty,\infty)$ loose starting bounds}
13:     for each feature $j\in[1,k]$ do
14:        $(s_{l},s_{r})\leftarrow\text{split}(s,j,D)$ create children}
15:        $p_{l}^{j}\leftarrow\text{FIND\_OR\_CREATE\_NODE}(G,s_{l},{\color[rgb]{1,0,0}d_{l}},{\color[rgb]{1,0,0}d},{\color[rgb]{1,0,0}d^{\prime}+1})$
16:        $p_{r}^{j}\leftarrow\text{FIND\_OR\_CREATE\_NODE}(G,s_{r},{\color[rgb]{1,0,0}d_{l}},{\color[rgb]{1,0,0}d},{\color[rgb]{1,0,0}d^{\prime}+1})$
@@
17:        $^{\prime}\leftarrow\min(lb^{\prime},p_{l}^{j}.\text{lb}+p_{r}^{j}.\text{lb})$ lb ) { create bounds as if $j$ were chosen for splitting}
18:        $^{\prime}\leftarrow\min(ub^{\prime},p_{l}^{j}.\text{ub}+p_{r}^{j}.\text{ub})$ ub )
19:     end for
20:     if $p.\text{lb}\neq lb^{\prime}$ or $p.\text{ub}\neq ub^{\prime}$ then { signal the parents if an update occurred}
21:        $p.\text{ub}\leftarrow\min(p.\text{ub},ub^{\prime})$
22:        $p.\text{lb}\leftarrow\min(p.\text{ub},\max(p.\text{lb},lb^{\prime}))$
23:        for $p_{\pi}\in G.\text{parent}(p)$ do { propagate information upwards}
24:           $Q.\text{push}((p_{\pi}.\text{id},{\color[rgb]{1,0,0}d^{\prime}-1}),\text{priority}=1)$
25:        end for
26:     end if
27:     if $p.\text{lb}\geq p.\text{ub}$ then
28:        continue { problem solved just now}
29:     end if
30:     if $d^{\prime}<d_{l}$ then
31:        for each feature $j\in[1,M]$ do { loop, enqueue all children}
32:           repeat line 14-16 { fetch $p_{l}^{j}$ and $p_{r}^{j}$ in case of update}
33:           $^{\prime}\leftarrow p_{l}^{j}.\text{lb}+p_{r}^{j}.\text{lb}$ lb
34:           $^{\prime}\leftarrow p_{l}^{j}.\text{ub}+p_{r}^{j}.\text{ub}$ ub
35:           if $lb^{\prime}<ub^{\prime}$ and $lb^{\prime}\leq p.\text{ub}$ then
36:              $Q.\text{push}((s_{l},{\color[rgb]{1,0,0}d+1}),\text{priority}=0)$
37:              $Q.\text{push}((s_{r},{\color[rgb]{1,0,0}d+1}),\text{priority}=0)$
38:           end if
39:        end for
40:     end if
41:  end while
42:  return $\mathcal{G}$
Algorithm 11 FIND_OR_CREATE_NODE($G$, $s$, $d_{l}$, $d$, $d^{\prime}$)
@@
0:  $G,s,d_{l},d,d^{\prime}$Graph, subproblem, lookahead depth, overall depth budget, current depth}
@@
4:     $p.\text{id}\leftarrow s$ identify $p$ by $s$}
@@
6:     $p.\text{ub},p.\text{lb}\leftarrow{\color[rgb]{1,0,0}\text{get\_bounds}(D(s),d_{l},d^{\prime},d)}$
@@
Algorithm 12 get_bounds($D$, $d_{l}$, $d^{\prime}$, $d$) $\to$ lb, ub
0:  $D$, $d_{l}$, $d^{\prime}$, $d$support, positive sample set, lookahead depth, current depth, overall depth budget}
@@
8:     $lb\leftarrow$ Equivalent points bound (Lin et al., 2020)
@@
Algorithm 13 extract__\__tree($D$,$\mathcal{G}$, $d_{l}$)
0:  $D$,$\mathcal{G}$, $d_{l}$Dataset, Dependency graph of search space, lookahead depth}
@@
2:  $t\leftarrow$ (Leaf predicting the majority label in $D$)
3:  $ub\leftarrow\lambda+$ (proportion of $D$ that has the minority label)
@@
13:     $t_{left}=$ extract__\__tree($D(f_{opt})$, $\mathcal{G}(f_{opt})$, $d_{l}-1$)
14:     $t_{right}=$ extract__\__tree($D(\bar{f_{opt}})$, $\mathcal{G}(\bar{f_{opt}})$, $d_{l}-1$)
@@
Algorithm 14 $\ell$, D, $\lambda$, $d_{l}$, $d$)
0:  $\ell$, $D$, $\lambda$, $d_{l}$, $d$loss function, samples, regularizer, lookahead depth, depth budget}
1:  $\mathcal{G}$ = $\ell$, D, $\lambda$, $d_{l}$, $d$)
@@ A.13 Additional Experimental Results
Figure 20: Tradeoff between train loss and runtime for all algorithms tested, for different sparsity levels (measured by ##\## of leaves in the tree). Depth Budget $=5$.
Figure 21: Tradeoff between test loss and runtime for all algorithms tested, for different sparsity levels (measured by ##\## of leaves in the tree). Depth Budget $=5$.
Figure 22: Regularized training objective vs. training time (in seconds) for GOSDT vs. our algorithms. The size of the points indicates the number of leaves in the resulting tree. Both SPLIT and LicketySPLIT are much faster for most values of sparsity penalty $\lambda$, with the only potential slowdown being in the sub-second regime due to overhead costs. Depth Budget $=5$.
@@
Dataset Binarization Time (s) Algorithm Runtimes (s) Test Loss
@@
Table 10: Results (##\## leaves between 3–6). The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time.
Dataset Binarization Time (s) Algorithm Runtimes (s) Test Loss
@@
Table 11: Results (##\## leaves between 7–10). The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time.
Dataset Binarization Time (s) Algorithm Runtimes (s) Test Loss
@@
Table 12: Results (##\## leaves between 11–14). The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time.
Dataset Binarization Time (s) Algorithm # Leaves (95% CI) Runtimes (s) Test Loss
@@
Table 13: Comparing CART and SPLIT/LicketySPLIT for non sparse trees. The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time. We report the best tree with between $15-30$ leaves found during hyperparameter search.
@@
Table 14: SPLIT/LicketySPLIT for non-sparse trees. For this variant, we set $\lambda=1e-5$ and ran our algorithms over $5$ trials. We show the $95\%$ confidence interval. This shows that our algorithms are capable of producing non-sparse trees. We may not prefer to do this in practice if there are interpretability constraints or if we can get a well performing model with much fewer than $20$ leaves.

### v3

@@ A.9 Greedy Algorithm
[…]
@@ A.10 RESPLIT Algorithm
Algorithm 5 RESPLIT($\ell$, D, $\lambda$, $_{l}, dd$)
[…]
1: ModifiedTreeFARMS = TreeFARMS reconfigured to use get_\_bounds (Algorithm 1) whenever it encounters a new subproblem
[…]
@@
[…]
@@
13:  $t_{lookahead}=$ Enumerate_\_TreeFARMS_\_subtrees {For each node in this prefix tree, store the number of subtrees we can generate rooted at that node. This speeds up indexing}
@@
15:return tftf {Return in-place edited ModifiedTreeFARMS object}
@@ A.11 Indexing Trees in RESPLIT
  • Once the cumulative count is known, we start looping over the entire Rashomon set. For the $i^{th}$ index, we first obtain the corresponding prefix tree and then find the relative index of the $i^{th}$ tree within this prefix tree structure. For example, if we query the $500^{th}$ tree and our prefix contains trees indexed 400600400-600 in the Rashomon set, the relative index of the query tree within this prefix is $100$.
  • Using Algorithm 8, we proceed to recursively locate the relevant subtrees beyond the prefix. In particular, at a given node in the prefix, we have access to the number of sub-trees that can be formed with its left and right children. We use this information to create two separate indexes for the left and right child (seen in lines 9109-10)
@@
Algorithm 6 Enumerate_\_TreeFARMS_\_subtrees
[…]
@@
4:  Return lenlen($t_{lookahead}$), $t_{lookahead}$
@@
6: left_\_expansions, left_\_subtree == enumerate_\_treefarms_subtrees($t_{lookahead}$.left_\_child)
[…]
9: right_\_expansions, right_\_subtree == enumerate_\_treefarms_\_subtrees($t_{lookahead}$.right_\_child)
[…]
12:Return left_\_expansions $\times$ right_\_expansions, $t_{lookahead}$ {Total number of subtrees = cross product of left and right subtree count}
Algorithm 7 RESPLIT_\_Rset_\_Count(RESPLIT_\_obj)
0: RESPLIT_\_obj {The RESPLIT object output by Algorithm 5}
[…]
3:for $t_{lookahead}\in$ RESPLIT_\_obj do
@@
7:   $s_{count}=len(tf_{u})$ {Number of subtrees found for subproblem D(u)D(u)}
@@
Algorithm 8 get_leaf_subtree_at_idx($t_{lookahead}$, tree_\_idx)
[…]
@@
[…]
7: left_\_count $t_{lookahead}$.left_\_child.subtree_\_count {The number of subtrees that can be found rooted at this node}
8: right_\_count $t_{lookahead}$.right_child.subtree_\_count
9: right_\_idx == tree_\_idx % right_\_count
10: left_\_idx == tree_\_idx // right_\_count
11:treetree.left_\_child == get_leaf_subtree_at_idx($t_{lookahead}$.left_child.node, left_\_idx)
12:treetree.right_\_child == get_leaf_subtree_at_idx($t_{lookahead}$.right_child.node, right_\_idx)
13:return treetree
Algorithm 9 RESPLIT_\_indexing
@@
1: hash = $\emptyset$ {Dictionary to map global tree indices to tree objects}
@@
6:   start = p_counts[i1i-1]+1 {Start index for prefix ii}
@@
8:  end = p_counts[$i$] {End index for prefix ii}
@@ A.12 Modifications to Existing GOSDT / TreeFARMS Code
In this section, we detail the main modifications we made to the existing GOSDT and TreeFARMS codebase in order to set up SPLIT, LicketySPLIT, and RESPLIT. The algorithm components in red are the modifications - note that GOSDT and TreeFARMS both call these functions. TreeFARMS does some additional post-processing of the search trie to find the set of near-optimal trees - the details can be seen in (xin2022treefarms).
Algorithm 10 find_\_lookahead_\_tree($\ell$, $D$, $\lambda$, $_{l}, dd$)
[…]
6:$N=|D|$ {global dataset size}
7:while $p_{0}.\text{lb}\neq p_{0}.\text{ub}$ do
8:  $s,{\color[rgb]{1,0,0}d}^{\prime}\leftarrow Q.\text{pop}()$ {index of problem to work on}
9:  $p\leftarrow G.\text{find}(s)$ { find problem to work on}
10:  if $p.\text{lb}=p.\text{ub}$ then
11:   continue { problem already solved}
12:  end if
13:  $(lb^{\prime},ub^{\prime})\leftarrow(\infty,\infty)$ { loose starting bounds}
14:  for each feature $j\in[1,k]$ do
15:   $(s_{l},s_{r})\leftarrow\text{split}(s,j,D)$ { create children}
16:   $p_{l}^{j}\leftarrow\text{FIND\_OR\_CREATE\_NODE}(G,s_{l},{\color[rgb]{1,0,0}d_{l}},{\color[rgb]{1,0,0}d},{\color[rgb]{1,0,0}d^{\prime}+1},N)$
17:   $p_{r}^{j}\leftarrow\text{FIND\_OR\_CREATE\_NODE}(G,s_{r},{\color[rgb]{1,0,0}d_{l}},{\color[rgb]{1,0,0}d},{\color[rgb]{1,0,0}d^{\prime}+1},N)$
@@
18:   $lb^{\prime}\leftarrow\min(lb^{\prime},p_{l}^{j}.\text{lb}+p_{r}^{j}.\text{lb})$ { create bounds as if $j$ were chosen for splitting}
19:   $ub^{\prime}\leftarrow\min(ub^{\prime},p_{l}^{j}.\text{ub}+p_{r}^{j}.\text{ub})$
20:  end for
21:  if $p.\text{lb}\neq lb^{\prime}$ or $p.\text{ub}\neq ub^{\prime}$ then { signal the parents if an update occurred}
22:   $p.\text{ub}\leftarrow\min(p.\text{ub},ub^{\prime})$
23:   $p.\text{lb}\leftarrow\min(p.\text{ub},\max(p.\text{lb},lb^{\prime}))$
24:   for $p_{\pi}\in G.\text{parent}(p)$ do { propagate information upwards}
25:    $Q.\text{push}((p_{\pi}.\text{id},{\color[rgb]{1,0,0}d^{\prime}-1}),\text{priority}=1)$
26:   end for
27:  end if
28:  if $p.\text{lb}\geq p.\text{ub}$ then
29:   continue { problem solved just now}
30:  end if
31:  if $d^{\prime}<d_{l}$ then
32:   for each feature $j\in[1,M]$ do { loop, enqueue all children}
33:    repeat line 14-16 { fetch $p_{l}^{j}$ and $p_{r}^{j}$ in case of update}
34:    $lb^{\prime}\leftarrow p_{l}^{j}.\text{lb}+p_{r}^{j}.\text{lb}$
35:    $ub^{\prime}\leftarrow p_{l}^{j}.\text{ub}+p_{r}^{j}.\text{ub}$
36:    if $lb^{\prime}<ub^{\prime}$ and $lb^{\prime}\leq p.\text{ub}$ then
37:     $Q.\text{push}((s_{l},{\color[rgb]{1,0,0}d+1}),\text{priority}=0)$
38:     $Q.\text{push}((s_{r},{\color[rgb]{1,0,0}d+1}),\text{priority}=0)$
39:    end if
40:   end for
41:  end if
42:end while
43:return $\mathcal{G}$
Algorithm 11 FIND_OR_CREATE_NODE($G$, $s$, $_{l}, dd$, $d^{\prime}$, N)
@@
0:$G,s$, $d_{l},d,d^{\prime}$, NN {Graph, subproblem, lookahead depth, overall depth budget, current depth, global dataset size}
@@
4:  $p.\text{id}\leftarrow s$ { identify $p$ by ss}
@@
6:  $p.\text{ub},p.\text{lb}\leftarrow{\color[rgb]{1,0,0}\text{get\_bounds}({\color[rgb]{0,0,0}D(s)},d_{l},d^{\prime},d},N{\color[rgb]{1,0,0})}$
@@
Algorithm 12 get_bounds($D$, $_{l}, dd^{\prime}$, $d$, NN) $\to$ lb, ub
0:$D$, $_{l}, dd^{\prime}$, $d$, NN {support, lookahead depth, current depth, overall depth budget, global dataset size}
@@
8:  $lb\leftarrow$ Equivalent points bound (gosdt)
@@
Algorithm 13 extract_\_tree($D$, $\mathcal{G}$, $d_{l}$)
[…]
@@
[…]
@@
[…]
@@
Algorithm 14 ModifiedGOSDT($\ell$, D, $\lambda$, $_{l}, dd$)
[…]
@@ A.13 Additional Experimental Results
Figure 20: Tradeoff between train loss and runtime for all algorithms tested, for different sparsity levels (measured by #\# of leaves in the tree). Depth Budget =5=5.
Figure 21: Tradeoff between test loss and runtime for all algorithms tested, for different sparsity levels (measured by #\# of leaves in the tree). Depth Budget =5=5.
Figure 22: Regularized training objective vs. training time (in seconds) for GOSDT vs. our algorithms. The size of the points indicates the number of leaves in the resulting tree. Both SPLIT and LicketySPLIT are much faster for most values of sparsity penalty $\lambda$, with the only potential slowdown being in the sub-second regime due to overhead costs. Depth Budget =5=5.
@@
Dataset
Binarization
Time (s)
Algorithm Runtimes (s) Test Loss
@@
Table 10: Results (#\# leaves between 3–6). The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time.
Dataset
Binarization
Time (s)
Algorithm Runtimes (s) Test Loss
@@
Table 11: Results (#\# leaves between 7–10). The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time.
Dataset
Binarization
Time (s)
Algorithm Runtimes (s) Test Loss
@@
Table 12: Results (#\# leaves between 11–14). The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time.
Dataset
Binarization
Time (s)
Algorithm
# Leaves
(95% CI)
Runtimes (s) Test Loss
@@
Table 13: Comparing CART and SPLIT/LicketySPLIT for non sparse trees. The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time. We report the best tree with between 153015-30 leaves found during hyperparameter search.
@@
Table 14: SPLIT/LicketySPLIT for non-sparse trees. For this variant, we set $\lambda=1e-5$ and ran our algorithms over 55 trials. We show the $95\%$ confidence interval. This shows that our algorithms are capable of producing non-sparse trees. We may not prefer to do this in practice if there are interpretability constraints or if we can get a well performing model with much fewer than 2020 leaves.